OpenStudy (sidsiddhartha):

another one :( a series of chords is drawn to the ellipse \[\large \frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] so that their projections on a straighe line which is inclined at an angle \[\large \alpha \] to the major axis are all of constant length C. find the locus of their middle point

OpenStudy (sidsiddhartha):

 here ive taken a point (h,k) so the equation of the chord will be \[\frac{ x-h }{ \cos \theta }=\frac{ y-k }{ \sin \theta }=r \\whuch ~gives~me \\x=h+rcos \theta \\y=k+rsin \theta\]